OpenStudy (anonymous):

Robin used to earn $25/h. Now she earns 1/6 more money per hour. Which expressions represent how much money Robin earns per hour now?

OpenStudy (texaschic101):

25 + 1/6(25) = 25 + 25/6 = 150/6 + 25/6 = 175/6 = 29 1/6 or 29.1666 rounds to $ 29.17 per hr.

OpenStudy (anonymous):

Um.Okay.I wanted you to explain it.But THANK YOU!If you ever answer one of my question just know PLEASE EXPLAIN IT!